OpenStudy (anonymous):

Solve the rational equation: 1/3 + 1/3 = 1/x Please explain so I can understand how to do it...

OpenStudy (anonymous):

Well, from my veiw, it appear as though one third and one third equal 1 something.

OpenStudy (anonymous):

First you simplify: 1/3 + 1/3 = (1+1)/3 = 2/3 Then you examine: 2/3 = 1/x x = ? Next you multiply the left side only by 0.5/0.5 (which equals 1): (2/3)*(0.5/0.5) = 1/1.5 => x = 1.5

OpenStudy (anonymous):

1/3+1/3 [here since the denominator is same you can write this as (1+1)/3] so u get 2/3=1/x first multiply both sides of the equation by x and after that both sides of the equation by 3. then keep 'x' on one side and the rest terms on the other side of the '=' sign. u get your answer...